[ the length and width of a rectangle are 6m and 4m. when each dimension is increased by the same amount, the area of the new RECTANGLE is 50 squared meters. find the dimensions of the new rectangle to the nearest tenth of a meter. ]
Answer by TimothyLamb(4379)

original rectangle:
a = 6 * 4 = 24
---
new rectangle:
A = (6 + n)(4 + n) = 50
(6 + n)(4 + n) = 50
24 + 10n + nn = 50
nn + 10n - 26 = 0
---
the above quadratic equation is in standard form, with a=1, b=10, and c=-26
---
to solve the quadratic equation, plug this:
1 10 -26
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
2.14142843
-12.1414284
---
the negative root doesn't make sense for a length, so use the positive root:
n = 2.14142843 meters
---
Answer:
new rectangle dimensions:
a = 6.1 m
b = 8.1 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php